OpenStudy (anonymous):

Please check my answer Complete the proof Given: 3x+2y=12 Prove: 3=y 3x+2y=12;x=2- Given 3(2)+2y=12- Substitution property of equality 2y=6- Multiplication property of equality y=3- Division property of equality 3=y- Reflexive property

OpenStudy (anonymous):

Whats the problem ?:) you can check for your self so easy ! just get : x=2 y=3 then we have : 3*2+2*3=12 ? ;)

OpenStudy (anonymous):

caz your equation was : 3x+2y=12 3*(2)+2*(3)=12 right or not ?!:)

OpenStudy (anonymous):

i need to know if my reasons are right

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

if you want to know it s right or not get your numbers in x and y or ...